Solving systems of three equations:
3x + 4y + z = 14
2y + 7z = 25
-2y + 3z = 5

Answer by checkley77(12844)

3x + 4y + z = 14
2y + 7z = 25
-2y + 3z = 5 ADD.
-------------------
10Z=30
Z=30/10
Z=3 ANS
2Y+7*3=25
2Y+21=25
2Y=25-21
2Y=4
Y=4/2
Y=2 ANS.
3X+4*2+3=14
3X+8+3=14
3X=14-8-3
3X=3
X=3/3
X=1 ANS.